An airplane travels 1800km in 6 hours flying with the wind. It travels only two thirds as far in the same time against the wind. Find the speed of the wind. 
Answer by Alan3354(69443)

An airplane travels 1800km in 6 hours flying with the wind. It travels only two thirds as far in the same time against the wind. Find the speed of the wind.
-----------
1800 km/6 hr = 300 km/hr
2/3 of that = 200 km/hr
-----------
wind speed = 1/2 the difference = 50 km/hr